OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (anonymous):

To start, you know that it must be geometric, meaning that the answer can't be A or B. Second, a3/a2 = a2/a1 = -4. Since the ratio between one to the next is -4, you know it can't be C, so the answer is D D.